# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 7.10. Regionwise seeds
    $c_{*}:=\min_{\partial\Omega}\big(H(g^{\rm in})-|\mbox{tr}_{\partial\Omega}k^{\rm in}|\big)>0,\qquad\text{Rad}(\Omega)=R_{*}<\frac{3\pi}{2\,c_{*}}.$
  2. (ii)
    Interface geometry. On the interface sphere $S_{-a,\underline{u}_{0}}=\partial\Omega\cap\partial\mathcal{M}_{2}$, set
    $\big(H(g^{\rm in}),\ \mbox{tr}_{S}k^{\rm in},\ \widehat{k}^{\rm in}\big)\big|_{S_{-a,\underline{u}_{0}}}=\big(H^{\rm dn},\ \mbox{tr}_{S}k^{\rm dn},\ \widehat{k}^{\rm dn}\big),$
    where the right side is imported from the double–null development (including the incoming shear $\widehat{\underline{\chi}}$ through the standard identification with tangential pieces of $k$).
No constraint is imposed yet in the interior of $\Omega$.
(II) Characteristic seed on $\mathcal{M}_{2}$ (from the double–null slab).
On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$ take the induced Cauchy data $(g^{\rm dn},k^{\rm dn})$ coming from the slab solution with scale–critical control on $(\widehat{\chi},\mbox{tr}\chi,\widehat{\underline{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\omega,\underline{\omega})$ and the shear budget $\int|\widehat{}\underline{\chi}_{0}|^{2}$ as prescribed in the characteristic problem. In particular, the transport system for $\widehat{\underline{\chi}}$ and $\mbox{tr}\underline{\chi}$ along $e_{4}$ is part of this data, hence the incoming focusing appears explicitly on $S_{-a,\underline{u}_{0}}$.
(III) Kerr seed on $\mathcal{M}_{3}$.
On the AF exterior choose a fixed Kerr slice $(g^{K}_{m,a},\pi^{K}_{m,a})$ with mass-angular momentum parameters $(m,J)$ and arbitrary Poincaré charges $(E,P,J,C)$ to be tuned later. Denote this by $(g^{\rm ex},k^{\rm ex})$.
7.11. Partition of unity and raw splice
Choose nested collars and annuli:
$\mathcal{C}_{1}\Subset\mathcal{M}_{1},\quad\mathcal{C}_{2}\Subset\mathcal{M}_{2},\quad A_{R}:=\{R<r<2R\}\Subset\mathcal{M}_{3},$
so that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are disjoint, and $A_{R}$ lies far out. Let $\{\xi_{1},\xi_{2},\xi_{3}\}$ be a smooth partition subordinate to $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$, with $\nabla\xi_{i}$ supported inside the chosen collars and annuli.
Define the splice (piecewise smooth) pair
$(g_{0},\pi_{0}):=\big(\xi_{1}g^{\rm in}+\xi_{2}g^{\rm dn}+\xi_{3}g^{\rm ex},\ \xi_{1}k^{\rm in}+\xi_{2}k^{\rm dn}+\xi_{3}k^{\rm ex}\big).$
By construction, $(g_{0},\pi_{0})$ agrees exactly with each seed away from the collars, and is smooth globally. Its constraint residual
$\mathcal{R}_{0}:=\Phi(g_{0},\pi_{0})$
is supported in $\operatorname{supp}\nabla\xi_{1}\cup\operatorname{supp}\nabla\xi_{2}\cup\operatorname{supp}\nabla\xi_{3}\subset\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup A_{R}$.
7.12. Compactly supported constraint correction (interior and interface)
Step A: kill $\mathcal{R}_{0}$ in $\mathcal{C}_{1}$ (purely interior).
Solve the projected adjoint equation on $\mathcal{C}_{1}$:
(199) $D\Phi^{*}_{(g_{0},\pi_{0})}(u_{1},X_{1})=\mathcal{T}_{1}(\mathcal{R}_{0})\quad\text{on }\mathcal{C}_{1},\qquad(u_{1},X_{1})\text{ compactly supported\ in }\mathcal{C}_{1},$
where $\mathcal{T}_{1}$ is a right–inverse onto the orthogonal complement of the local obstruction space. Solve the associated quadratic corrected system to get a compactly supported perturbation
$(\delta g_{1},\delta\pi_{1})=\mathcal{Q}_{1}(u_{1},X_{1})$
with $\Phi(g_{0}+\delta g_{1},\pi_{0}+\delta\pi_{1})=\mathcal{R}_{0}-\mathcal{P}_{1}(\mathcal{R}_{0})$ on $\mathcal{C}_{1}$, where $\mathcal{P}_{1}$ is the projection onto the local obstruction. Because $\mathcal{C}_{1}$ is strictly interior, no global charges are involved.
Step B: kill $\mathcal{R}_{0}$ in $\mathcal{C}_{2}$ (interface to double–null).
Repeat (199) on $\mathcal{C}_{2}$ to obtain a c.s. correction $(\delta g_{2},\delta\pi_{2})$ supported in $\mathcal{C}_{2}$, leaving the trace data on the common sphere $S_{-a,\underline{u}_{0}}$ unchanged. This guarantees that the double–null induced tangential geometry—hence the scalars built from $(\mbox{tr}\chi,\mbox{tr}\underline{\chi})$ and the shear $\widehat{\underline{\chi}}$—is preserved on the interface.
After Steps A and B we have
$\Phi(g_{1},\pi_{1})=\Phi\big((g_{0},\pi_{0})+(\delta g_{1}+\delta g_{2},\delta\pi_{1}+\delta\pi_{2})\big)=\mathcal{R}_{R}$
with $\operatorname{supp}\mathcal{R}_{R}\subset A_{R}$.
7.13. Exterior compactly supported correction with Kerr parameter tuning
On the large annulus $A_{R}$, perform the compactly supported exterior correction: solve
$D\Phi^{*}_{(g_{1},\pi_{1})}(u_{3},X_{3})=\mathcal{T}_{R}(\mathcal{R}_{R})\quad\text{on }A_{R},$
producing a compactly supported perturbation $(\delta g_{3},\delta\pi_{3})$ supported in $A_{R}$. The remaining obstruction lives in the global $10$-dimensional space $\mathcal{K}$; adjust the Kerr parameters $(E,P,J,C)$ so that the parameter map hits the obstruction and annihilates it. Concretely, choose $(E,P,J,C)$ so that
$\Pi_{\mathcal{K}}\big(\Phi(g_{1}+\delta g_{3},\pi_{1}+\delta\pi_{3})\big)=0,$
which is possible by the surjectivity argument. The outcome is a vacuum pair
$(g^{\sharp},\pi^{\sharp})$
with
$\Phi(g^{\sharp},\pi^{\sharp})=0,\qquad(g^{\sharp},\pi^{\sharp})=\begin{cases}(g^{\rm in},k^{\rm in})&\text{on }\mathcal{M}_{1}\setminus\mathcal{C}_{1},\\ (g^{\rm dn},k^{\rm dn})&\text{on }\mathcal{M}_{2}\setminus\mathcal{C}_{2},\\ (g^{K}_{m,a},\pi^{K}_{m,a})&\text{on }\{r\geq 2R\},\end{cases}$
and smooth across all interfaces.
7.14. Yau–radius prescription and MOTS/future AH behavior
By design, the boundary geometry on $\partial\Omega$ (of $\mathcal{M}_{1}$) is that of the double–null interface sphere, with
@@ 7.14. Yau–radius prescription and MOTS/future AH behavior
$c_{*}=\min_{\partial\Omega}\big(H-|\mbox{tr}_{\partial\Omega}k|\big)>0,\qquad\text{Rad}(\Omega)=R_{*}<\frac{3\pi}{2\,c_{*}}.$
The compactly supported corrections never alter $\partial\Omega$, hence the Yau barrier conditions and the prescribed Yau radius are retained on the initial slice. Consequently, there is no MOTS on t=at=-a. On the other hand, the shear content imported on $\mathcal{M}_{2}$ (including $\widehat{\underline{\chi}}$) guarantees MOTS formation in the future development (as per the scale-critical characteristic criteria) while the initial slice remains regular.
7.15. Local evolution in harmonic gauge
@@
We now solve the vacuum Einstein equations
$\mathrm{Ric}[\widehat{g}]=0$
in the harmonic gauge
(200) $\Gamma^{\mu}:=\widehat{g}^{\alpha\beta}\Gamma[\widehat{g}]^{\mu}_{\alpha\beta}=0.$
Under (200) the field equations reduce to the quasilinear wave system

### v4

@@ 6.4. Interior data on $\mathcal{M}_{1}$: strict Yau–radius subcriticality with boundary control
$\min_{\partial\Omega}\Big(H(\tilde{g})-|\mbox{tr}_{\partial\Omega}\tilde{k}|\Big)\geq c_{*}>0,\qquad\text{Rad}(\Omega)\leq R_{*}<R_{\mathrm{crit}}(c_{*}),$
and therefore $\mathcal{M}_{1}$ contains no MOTS on the initial slice t=at=-a. Since all subsequent smoothing and interface adjustments are performed by compactly supported deformations away from $\partial\Omega$ with size $O(a^{-3/2})$ in uniformly–local norms, both the boundary gap and the radius bound are stable under these operations.
It remains to ensure that the same nonexistence property holds on the full initial slice $\mathcal{M}_{-a}$. This is provided by the following proposition, which shows that the Schoen–Yau barrier–radius inequality persists across the glued regions and excludes trapped or marginally outer trapped surfaces on $\mathcal{M}_{-a}$.
Proposition 6.1 (Absence of MOTS on the initial slice).
Let $\mathcal{M}_{-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$ be the initial Cauchy slice constructed as follows: (i) $\mathcal{M}_{1}$ is the interior region with boundary sphere $S_{-a,0}=\partial\mathcal{M}_{1}$ carrying the prescribed boundary geometry from the double–null interface; (ii) $\mathcal{M}_{2}=\mathcal{M}_{-a}\cap D_{a,1}$ is the characteristic gluing region with induced data from the characteristic development $D_{a,1}$; (iii) $\mathcal{M}_{3}=\mathcal{M}_{ext}$ is an exterior Kerr slice (in the domain of outer communication). Assume that the boundary data on $S_{-a,0}$ satisfy the strict Yau–barrier inequality in the form of condition (c) of Theorem 2.2, and that $\text{Rad}(\mathcal{M}_{1})$ is chosen by
(237) $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1}),~a\gg 1$
Then $\mathcal{M}_{-a}$ contains no closed marginally outer trapped surface and no closed trapped surface.
Proof.
With the choice (237), a direct expansion gives
(238) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}-\frac{2}{a}=\frac{2}{a^{2}}+O(a^{-3}),$
and likewise
(239) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})\sqrt{1+\frac{1}{10a}}}-\frac{2}{a}=\frac{8}{5a^{2}}+O(a^{-3}).$
Under condition (c) of Theorem 2.2, i.e.
(240) $\frac{17}{10a^{2}}<\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}||\underline{\hat{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\underline{\hat{\chi}}|^{2}(u^{\prime\prime},\epsilon)\,du^{\prime\prime}\,du^{\prime}<\frac{19}{10a^{2}},$
the computation carried out in the proof of Theorem 2.2 yields the quantitative defect
(241) $\Big(H-|\kappa|\Big)\big|_{S_{-a,0}}-\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}\leq-\frac{1}{10a^{2}}+O(a^{-5/2}).$
In particular, for $a\gg 1$ the right-hand side is strictly negative. Hence $S_{-a,0}$ violates the Yau criterion (cf. [44] and the formulation 2.1) with a strict margin. By the Yau barrier theorem, $\mathcal{M}_{1}$ contains no closed MOTS (and thus no closed trapped surface).
By construction, $\mathcal{M}_{3}$ is a Kerr slice lying in the domain of outer communication; in particular it contains no closed trapped surface and no closed MOTS.
Let $S\subset\mathcal{M}_{2}$ be any embedded $2$-sphere arising as a section of the outgoing null foliation in $D_{a,1}$. Along each incoming null generator, $\mbox{tr}\chi$ satisfies the Raychaudhuri transport equation
(242) $\nabla_{3}\mbox{tr}\chi+\frac{1}{2}\mbox{tr}\underline{\chi}\,\mbox{tr}\chi=2\underline{\omega}\,\mbox{tr}\chi+2\mbox{div }\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\underline{\hat{\chi}}$.
Using the bootstrap bounds available in $D_{a,1}$ (in particular the standard estimates on $\underline{\omega},\eta,\rho,\hat{\chi},\underline{\hat{\chi}}$ inherited from the characteristic construction) and Grönwall along the $\nabla_{3}$–flow, one obtains the quantitative perturbative estimate
(243) $\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-1/2}|u|^{-1}\qquad\text{throughout }\mathcal{M}_{2}.$
On $\mathcal{M}_{2}$ one has $|u|\sim a$ (since $\mathcal{M}_{2}=\mathcal{M}_{-a}\cap D_{a,1}$ is localized near the interface), hence (243) improves to
(244) $\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-3/2},\qquad\text{on }\mathcal{M}_{2}.$
Therefore, for $a\gg 1$, $0>u\in[u_{\infty},-a]$
@@
$\mbox{tr}\chi\geq\frac{2}{|u|}-Ca^{-3/2}\geq ca^{-1}>0\qquad\text{on every section }S\subset\mathcal{M}_{2},$
with $c>0$ universal. In particular, no $S\subset\mathcal{M}_{2}$ can satisfy $\mbox{tr}\chi=0$, so $\mathcal{M}_{2}$ contains no closed MOTS; moreover, $\mbox{tr}\chi$ cannot be negative on such a closed section, so there are no closed trapped surfaces in $\mathcal{M}_{2}$. This completes the proof. ∎
We provide the following proposition sketching the gluing mechanism to construct data on the slice $\mathcal{M}_{-a}$. This is a direct application of the Corvino-Schoen gluing technique [22]. This is slightly less important in the sense that we are ultimately interested in the interior dynamics and dynamical formation of MOTS. This study deserves a separate rigorous attention which we leave it for future work. But we provide a sketch of the proof for completeness.
Proposition 6.2 (Construction of constraint data with interior–characteristic–Kerr matching).
Fix an integer $N\geq 4$, $s\leq 3$ for suffiicnetly large $\mathbb{N}$, and $a\gg 1$. There exist a smooth Riemannian metric $\widetilde{g}$ and a smooth symmetric 22–tensor $k$ on the Cauchy slice
$\mathcal{M}_{-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext},$
@@
such that $(\widetilde{g},k)$ solve the vacuum Einstein constraint equations on $\mathcal{M}_{-a}$, i.e.